OpenStudy (anonymous):

please help me. simplest steps gets medal and a fan. 9.6+0.20x = 0.16(x+12) I got -192 the first time but i KNOW thats not right.

OpenStudy (anonymous):

hold on i erased it lready

OpenStudy (aravindg):

wait a sec

OpenStudy (anonymous):

it is -192

OpenStudy (aravindg):

your answer is right!!

OpenStudy (aravindg):

well done @Luckycutie1

OpenStudy (anonymous):

9.6+0.2(-192)=0.16(-192+12) 9.6-38.4=0.16(-180) -28.8=-28.8

OpenStudy (anonymous):

the answer isn't -192. its 24 it says so on the answer sheet. but i need the steps. i did this.. 9.6+0.20x =0.16(x+12) 9.6+0.20x = 0.16+1.92 7.68+0.20x = 0.16 7.68 = -0.04 -192??

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

I did the same thing

OpenStudy (anonymous):

-0.04x

OpenStudy (aravindg):

yep you are right !

OpenStudy (anonymous):

try to solve it with 24 and is not the same answer at the end

OpenStudy (anonymous):

9.6+0.20x=0.16(x+12) 9.6+0.20x=0.16x+1.92 9.6+0.04x=1.92 0.04x=-7.68 x=-192 9.6+0.20(-192)=0.16(-192+12) 9.6+(-38.4)=-28.8 -28.8=-28.8 sometimes answer sheets are wrong

OpenStudy (anonymous):

well i peeked at my teachers answer sheet and it 24 weird. do u want me to give you the full question